OpenStudy (anonymous):

What is the equation of a line passing through the point (a, b) and having a slope of b?

OpenStudy (imstuck):

Do this like you would if there were numbers instead of letters inside the parenthesis and as if there was a slope with a value instead of just b. Do you know the point slope equation?

OpenStudy (anonymous):

im in 7th grade doing 10th grade math and this is the pretest

OpenStudy (imstuck):

Ok, this is the point-slope form of an equation. Pretend that a is your x coordinate, x1, and b is your y coordinate, y1. ok?

OpenStudy (imstuck):

\[y-y _{1}=m(x-x _{1})\]

OpenStudy (imstuck):

y1 is b, x1 is a, and the slope (m) is b. So this is your equation in point-slope form:\[y-b=b(x-a)\]

OpenStudy (imstuck):

You could then expand that and put it into slope intercept form, y = mx + b, by moving the -b away from the y and distributing the b (slope) into the parenthesis. That would then give you this: y=bx-ab+b

OpenStudy (anonymous):

y = −bx − b(1 − a) y = bx − b(a − 1) y=1/bx+b[a-1] y=-1/bx+b[1-a] these are my answer choices

OpenStudy (anonymous):

y = −bx − b(1 − a) y = bx − b(a − 1) y=1/bx+b[a-1] y=-1/bx+b[1-a] these are my answer choices